Gladys leaves a town driving at a rate of 40 miles per hour. Two hours later, Reena leaves from the same places traveling the same route. she catches Gladys in 5 hours and 20 minutes. how fast was Reena traveling.
Answer by mananth(16946)

Gladys 40 mph
Reena x mph

Reena starts after 2hours
Catchup time = 5hours 20minutes = 7.33 hours

Time taken by gladys= 7.33 hours

Distance is same
7.33*40 = 5*x
293.33 =5*x
/5
x=58.67 mph Reena's speed